# Resilient Topology-Aware Coordination for Dynamic 3D UAV Networks under Node Failure

### v1

@@ 4.1.2. Global Energy Efficiency
(10) $E_{\text{eff}}(t)=\frac{R_{\text{sum}}(t)}{\sum_{k\in\mathcal{V}(t)\setminus{\text{GBS}}}(P_{k}^{\text{fly}}(t)+P_{\text{comm}})+P_{\text{GBS}}},$
where $P_{k}^{\text{fly}}(t)$ is the propulsion power required for UAV $k$ to maintain flight at its instantaneous speed, while $P_{\text{GBS}}$ represents the constant power consumption of the stationary ground infrastructure. By evaluating $E_{\text{eff}}(t)$, we can assess whether the observed network resilience is achieved through energy-efficient maneuvers or via aggressive, high-speed propulsion that significantly reduces the endurance of the swarm.
4.1.3. Coverage and QoS Satisfaction
@@
4.1.4. Fairness Metrics
We employ Jain’s Fairness Index across two dimensions to ensure equitable service. First, Rate Fairness $\mathcal{J}_{\text{R}}(t)$ ensures comparable data speeds among users:
(11) $\mathcal{J}_{\text{R}}(t)=\frac{(\sum_{m=1}^{M}R_{m}(t))^{2}}{M\sum_{m=1}^{M}(R_{m}(t))^{2}+\epsilon},$
where $\epsilon$ is a small constant for numerical stability. Second, Load Fairness $\mathcal{J}_{\text{L}}(t)$ prevents congestion by evaluating the distribution of connected users across the active UAV set $\mathcal{V}(t)\setminus\{\text{GBS}\}$.
4.1.5. Service Continuity and Handoff Cost
@@
4.1.6. Composite Utility Function
@@
(13) $U(t)=U_{\text{QoS}}(t)-\lambda_{\text{HO}}C_{\text{HO}}(t),$
@@
(14) $U_{\text{QoS}}(t)=\lambda_{\text{EE}}\tilde{E}_{\text{eff}}(t)+\lambda_{\text{JR}}\mathcal{J}_{\text{R}}(t)+\lambda_{\text{JL}}\mathcal{J}_{\text{L}}(t)+\lambda_{\text{Cov}}C_{\text{cov}}(t)+\lambda_{\text{Min}}\tilde{R}_{\min}(t).$
@@
4.2. Optimization Problem Statement and Feasibility Analysis
@@
We aim to optimize the joint 3D trajectories of all operational UAVs within the dynamic set $\mathcal{V}_{U}(t)=\mathcal{V}(t)\setminus\{\text{GBS}\}$ to maximize the long-term time-averaged utility. The optimization problem is formulated as follows:
(15a) $(\textbf{P1}):\max_{\mathbf{P}}\quad$ $\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}[U(t)]$
(15b) s.t. $\mathbf{p}_{k}(t+1)=\mathbf{p}_{k}(t)+\mathbf{v}_{k}(t)\Delta t,\quad\forall k\in\mathcal{V}_{U}(t)$
(15c) $|\mathbf{v}_{k}(t)|\leq V_{\max}^{\text{UAV}},\quad\forall k\in\mathcal{V}_{U}(t)$
(15d) $H_{\min}\leq z_{k}(t)\leq H_{\max},\quad\forall k\in\mathcal{V}_{U}(t)$
(15e) $\mathbf{p}_{k}(t)\in\mathcal{D},\quad\forall k\in\mathcal{V}_{U}(t)$
(15f) $|\mathbf{p}_{i}(t)-\mathbf{p}_{j}(t)|\geq d_{\text{safe}},\quad\forall i,j\in\mathcal{V}(t),i\neq j$
@@
The constraints specified in (15) define the physical boundaries of the heterogeneous network under potential topological disruptions. Specifically, the kinematic feasibility of the surviving UAV swarm is governed by (15b) and (15c). Flight corridor constraints in (15d) and (15e) restrict the UAV altitudes $z_{k}(t)$ and coordinates $\mathbf{p}_{k}(t)$ within compliant airspace. Finally, the collision avoidance requirement in (15f) mandates a minimum safety distance $d_{\text{safe}}$ between any pair of active nodes, which includes both UAV-to-UAV and UAV-to-GBS separations. This is critical for preventing secondary hardware failures during rapid reconfiguration when the remaining agents move aggressively to fill coverage gaps left by failed nodes.
Computational Challenges and Feasibility:
@@
The optimization problem (P1) presents significant challenges that render classical analytical frameworks impractical for real-time deployment. First, the utility $U(t)$ involves intricate dependencies on 3D coordinates, probabilistic channels, and dynamic interference, creating a highly non-convex landscape where gradient-based methods frequently converge to suboptimal local optima. Second, the joint state-action space suffers from the curse of dimensionality because the complexity grows exponentially with the number of users $M$ and operational UAVs in the surviving set $\mathcal{V}_{U}(t)$. Third, the high-mobility environment and sudden node failures imply that optimal trajectories derived for a specific snapshot quickly become obsolete. Classical iterative methods, such as Successive Convex Approximation (SCA), require significant re-computation time whenever the network topology undergoes a rupture, failing to satisfy the millisecond-level responsiveness required for 6G applications.
@@
Consequently, the necessity for a low-latency, inference-based control paradigm is evident. By reformulating (P1) as a Deep Reinforcement Learning (DRL) task, the computational burden is shifted to the offline training phase. During online execution, the proposed TAG-MAPPO framework provides near-instantaneous control actions through a single forward pass, ensuring feasible real-time topology reconfiguration and rapid recovery after node failures.
4.3. Reformulation as a Multi-Agent DEC-POMDP
@@
To address the intractability of (P1), we reformulate the control problem as a Multi-Agent Decentralized Partially Observable Markov Decision Process (DEC-POMDP), characterized by the tuple $\langle\mathcal{K},\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\Omega,\mathcal{O},\gamma\rangle$.
4.3.1. Global State Space $\mathcal{S}$
@@
The global state $s_{t}\in\mathcal{S}$ characterizes the complete configuration of the heterogeneous network at time $t$. To ensure that the centralized critic evaluates the impact of node failures during the training phase, the global state incorporates the survival status of each agent. The state vector is defined as follows:
(16) $s_{t}=\left\{\left\{\mathbf{p}_{k}(t),\mathbf{v}_{k}(t),\xi_{k}(t)\right\}_{k\in\mathcal{K}_{U}},\mathbf{p}_{\text{GBS}},\left\{\mathbf{q}_{m}(t),u_{m}(t)\right\}_{m\in\mathcal{M}}\right\},$
@@
where $\mathcal{K}_{U}$ denotes the set of $K_{U}=4$ mobile UAVs. The inclusion of $\xi_{k}(t)$ captures the operational status of each UAV node. By incorporating the fixed GBS position $\mathbf{p}_{\text{GBS}}$ and the dynamic user positions $\mathbf{q}_{m}(t)$, the critic can estimate the value function based on the complete spatial relationship between infrastructure, aerial nodes, and ground users.This comprehensive representation allows the model to capture the complex dependencies required for effective centralized training and decentralized execution.
4.3.2. Local Observation Space $\Omega$
@@
In decentralized execution, each operational UAV $k\in\mathcal{V}_{U}(t)$ makes control decisions based on its local observation $o_{k}(t)$. To achieve topology-invariant learning and enhance spatial awareness, absolute coordinates are transformed into a UAV-centric relative frame. The observation vector is defined as follows:
(17) $o_{k}(t)=\left[\mathbf{x}_{k}^{\text{self}},\mathbf{x}_{k}^{\text{GBS}},\mathbf{X}_{k}^{\text{neigh}},\mathbf{x}_{k}^{\text{user}}\right],$
where each component is formulated to capture specific environmental features:
@@
  • Self-State ($\mathbf{x}_{k}^{\text{self}}$): This component contains the kinematic information of the agent itself, defined as $\mathbf{x}_{k}^{\text{self}}=[z_{k}(t),\mathbf{v}_{k}(t)]$. The altitude $z_{k}(t)$ and velocity vector $\mathbf{v}_{k}(t)$ are normalized based on the safe flight corridor and the maximum speed $V_{\max}^{\text{UAV}}$ respectively.
@@
  • GBS Anchor ($\mathbf{x}_{k}^{\text{GBS}}$): To provide a stable global reference point, the agent perceives its relative displacement to the fixed ground station. It is expressed as $\mathbf{x}_{k}^{\text{GBS}}=\mathbf{p}_{\text{GBS}}-\mathbf{p}_{k}(t)$. This feature allows the swarm to maintain a consistent coordinate origin even after the failure of other aerial nodes.
@@
  • Neighborhood Topology ($\mathbf{X}_{k}^{\text{neigh}}$): This set consists of relative features from neighboring UAVs within the communication range $R_{\text{comm}}$. For each operational neighbor $j\in\mathcal{V}_{U}(t)\setminus\{k\}$, the feature includes the relative position $\Delta\mathbf{p}_{kj}(t)=\mathbf{p}_{j}(t)-\mathbf{p}_{k}(t)$ and the velocity $\mathbf{v}_{j}(t)$. The combined set $\mathbf{X}_{k}^{\text{neigh}}=\{[\Delta\mathbf{p}_{kj}(t),\mathbf{v}_{j}(t)]\}_{j\in\mathcal{V}_{U}(t)\setminus\{k\}}$ is processed by the attention mechanism to dynamically re-weight coordination following failures.
@@
  • User Distribution Features ($\mathbf{x}_{k}^{\text{user}}$): These features are perceived as localized user densities or aggregated centroids within the service footprint. They provide the necessary spatial cues for tracking high-mobility vehicular platoons and identifying areas with high service demand.
4.3.3. Action Space $\mathcal{A}$
@@
We define a discrete action space $\mathcal{A}$ where each action $a_{k}(t)\in\mathcal{A}$ corresponds to a specific target velocity vector. To simulate the physical inertia of rotary-wing UAVs and ensure smooth motion, the actual velocity vector $\mathbf{v}_{k}(t+1)$ is updated according to the following kinematic rule:
(18) $\mathbf{v}_{k}(t+1)=\beta\mathbf{v}_{k}(t)+(1-\beta)\mathbf{v}_{\text{target}}(a_{k}(t))+\boldsymbol{\epsilon},$
@@
where $\beta\in[0,1)$ is the inertia coefficient and $\boldsymbol{\epsilon}$ denotes the small mechanical noise vector. This formulation ensures that the resulting trajectories remain continuous while the norm of the target velocity vector is pre-scaled to satisfy the maximum flight speed constraint $\|\mathbf{v}_{k}(t+1)\|\leq V_{\max}^{\text{UAV}}$ for all operational UAVs in $\mathcal{V}_{U}(t)$.
4.3.4. Cooperative Reward Function $\mathcal{R}$
To foster full cooperation among the heterogeneous nodes, all agents share a common reward signal $r(t)$ based on the composite utility defined in Section 4.1.6. The reward at each time step is set as $r(t)=U(t)$, which directly incorporates the weighted metrics of global energy efficiency, rate fairness, load distribution, coverage ratio, and the stability penalty for handoffs. The primary objective of the swarm is to find an optimal joint policy $\boldsymbol{\pi}$ that maximizes the expected discounted return:
(19) $J(\boldsymbol{\pi})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(t)\right],$
@@ 4.3.4. Cooperative Reward Function $\mathcal{R}$
where $\gamma\in[0,1)$ is the discount factor. By maximizing this return, the surviving UAVs are incentivized to autonomously re-configure their spatial topology to compensate for service gaps while maintaining the system stability constraints specified in (13) and (14). This collective reward mechanism ensures that individual agent actions are aligned with the global network performance objectives under dynamic node failure conditions.
@@
5. Topology-Aware Graph-based Multi-Agent Reinforcement Learning (TAG-MAPPO)
@@
In this section, we propose the Topology-Aware Graph-based Multi-Agent Reinforcement Learning (TAG-MAPPO) framework. This framework is specifically engineered to surmount the challenges of high-mobility tracking and rapid topology deformation caused by both user movement and sudden node failures. To address the performance degradation typical of existing solutions in volatile environments, TAG-MAPPO adopts a Centralized Training with Decentralized Execution (CTDE) paradigm augmented by a specialized attention mechanism. Recent literature confirms that graph-based learning is uniquely suited for such dynamic wireless resource management tasks because of its inherent scalability and generalization capabilities (Dai et al., 2025).
@@
5.1. Overview of the TAG-MAPPO Architecture
The overall architectural flow of the proposed TAG-MAPPO framework is illustrated in Figure 2. The processing pipeline is strictly structured to differentiate between the online execution phase and the offline centralized training phase.
Fig 2. The schematic architecture of the proposed TAG-MAPPO framework: (1) Decentralized Execution (Actor): Each UAV agent utilizes a lightweight, shared MLP encoder to map local observations $o_{k}(t)$ to actions, ensuring real-time responsiveness. (2) Centralized Training (Critic): The critic leverages a Topology-Aware Graph Attention (TA-GAT) mechanism to estimate state values $V_{\phi}(s_{t})$. This module employs a decoupled dual-path strategy. A Random Observation Shuffling (ROS) operator is applied to the neighbor feature set $\{\mathbf{h}_{j}\}_{j\in\mathcal{N}_{k}(t)}$ to ensure permutation invariance. Simultaneously, a bold red skip connection preserves the ego-state information $\mathbf{h}_{k}$ to maintain operational consistency. This architecture enables the framework to perform precise relational reasoning and rapid spatial reconfiguration, which ensures robust coordination even under sudden node failures.

### v2

@@ IV-A2 Global Energy Efficiency
$E_{\text{ee}}(t)=\frac{R_{\text{sum}}(t)}{\sum_{k\in\mathcal{V}(t)\setminus{\text{GBS}}}(P_{k}^{\text{fly}}(t)+P_{\text{comm}})+P_{\text{GBS}}},$ (10)
where $P_{k}^{\text{fly}}(t)$ is the propulsion power required for UAV $k$ to maintain flight at its instantaneous speed, while $P_{\text{GBS}}$ represents the constant power consumption of the stationary ground infrastructure. By evaluating $E_{\text{ee}}(t)$, we can assess whether the observed network resilience is achieved through energy-efficient maneuvers or via aggressive, high-speed propulsion that significantly reduces the endurance of the swarm.
IV-A3 Coverage and QoS Satisfaction
@@
IV-A4 Fairness Metrics
We employ Jain’s Fairness Index to ensure equitable service among users. Rate Fairness $\mathcal{J}_{\text{R}}(t)$ ensures comparable data speeds to prevent specific users from suffering severe throughput degradation:
$\mathcal{J}_{\text{R}}(t)=\frac{(\sum_{m=1}^{M}R_{m}(t))^{2}}{M\sum_{m=1}^{M}(R_{m}(t))^{2}+\epsilon},$ (11)
where $\epsilon$ is a small constant for numerical stability.
IV-A5 Service Continuity and Handoff Cost
@@
IV-A6 Composite Utility Function
@@
$U(t)=U_{\text{QoS}}(t)-\lambda_{\text{ho}}C_{\text{ho}}(t),$ (13)
@@
$\begin{split}U_{\text{QoS}}(t)=&\lambda_{\text{ee}}\tilde{E}_{\text{ee}}(t)+\lambda_{\text{jr}}\mathcal{J}_{\text{r}}(t)+\lambda_{\text{cov}}C_{\text{cov}}(t)\\ &+\lambda_{\text{min}}\tilde{R}_{\text{min}}(t).\end{split}$ (14)
@@
IV-B Optimization Problem Statement and Feasibility Analysis
@@
We aim to optimize the joint 3D trajectories of all operational UAVs in $\mathcal{V}_{U}(t)=\mathcal{V}(t)\setminus\{\text{GBS}\}$ to maximize long-term time-averaged utility. The optimization problem is:
$(\textbf{P1}):\max_{\mathbf{P}}\quad$ $\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}[U(t)]$ (15a)
s.t. $\mathbf{p}_{k}(t+1)=\mathbf{p}_{k}(t)+\mathbf{v}_{k}(t)\Delta t,\quad\forall k\in\mathcal{V}_{U}(t)$ (15b)
$|\mathbf{v}_{k}(t)|\leq V_{\max}^{\text{UAV}},\quad\forall k\in\mathcal{V}_{U}(t)$ (15c)
$H_{\min}\leq z_{k}(t)\leq H_{\max},\quad\forall k\in\mathcal{V}_{U}(t)$ (15d)
$\mathbf{p}_{k}(t)\in\mathcal{D},\quad\forall k\in\mathcal{V}_{U}(t)$ (15e)
$|\mathbf{p}_{i}(t)-\mathbf{p}_{j}(t)|\geq d_{\text{safe}},\quad\forall i,j\in\mathcal{V}(t),i\neq j$ (15f)
@@
The constraints in (15) define the physical boundaries under potential topological disruptions. Kinematic feasibility is governed by (15b) and (15c). Flight corridor constraints (15d) and (15e) restrict UAV altitudes $z_{k}(t)$ and coordinates $\mathbf{p}_{k}(t)$ within compliant airspace. Collision avoidance in (15f) ensures a minimum safety distance $d_{\text{safe}}$ between active nodes, including UAV-to-UAV and UAV-to-GBS separations. This is critical to prevent secondary failures during rapid reconfiguration when agents move aggressively to fill coverage gaps.
@@
Computational Challenges and Feasibility: Problem (P1) presents challenges that render classical analytical frameworks impractical for real-time deployment. First, utility $U(t)$ involves intricate dependencies on coordinates, probabilistic channels, and dynamic interference, creating a non-convex landscape where gradient-based methods frequently converge to local optima. Second, the joint state-action space is subject to the curse of dimensionality, as complexity grows exponentially with users $M$ and surviving UAVs $\mathcal{V}_{U}(t)$. Third, high-mobility and node failures mean optimal trajectories derived for specific snapshots quickly become obsolete. Classical iterative methods, such as Successive Convex Approximation (SCA), require significant re-computation time upon topological ruptures, failing the millisecond-level responsiveness required for 6G applications.
@@
Consequently, a low-latency, inference-based control paradigm is essential. By reformulating (P1) as a Deep Reinforcement Learning task, the computational burden shifts to offline training. During execution, the TAG-MAPPO framework provides near-instantaneous control through a single forward pass, ensuring real-time topology reconfiguration and rapid recovery.
IV-C Reformulation as a Multi-Agent DEC-POMDP
@@
To address the intractability of (P1), we reformulate the control problem as a Multi-Agent Decentralized Partially Observable Markov Decision Process (DEC-POMDP) defined by the tuple $\langle\mathcal{K},\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\Omega,\mathcal{O},\gamma\rangle$.
IV-C1 Global State Space $\mathcal{S}$
@@
The global state $s_{t}\in\mathcal{S}$ characterizes the complete network configuration at time $t$. To ensure the centralized critic evaluates node failures during training, the global state incorporates survival statuses. The state vector is:
$s_{t}=\left\{\left\{\mathbf{p}_{k}(t),\mathbf{v}_{k}(t),\xi_{k}(t)\right\}_{k\in\mathcal{K}_{U}},\mathbf{p}_{\text{GBS}},\left\{\mathbf{q}_{m}(t),u_{m}(t)\right\}_{m\in\mathcal{M}}\right\},$ (16)
@@
where $\xi_{k}(t)$ captures the operational status of each UAV $k\in\mathcal{K}_{U}$. By incorporating the GBS position $\mathbf{p}_{\text{GBS}}$ and dynamic user positions $\mathbf{q}_{m}(t)$, the critic estimates the value function based on spatial relationships between infrastructure, aerial nodes, and users. This representation allows the model to capture complex dependencies for centralized training and decentralized execution.
IV-C2 Local Observation Space $\Omega$
@@
In decentralized execution, each operational UAV $k\in\mathcal{V}_{U}(t)$ makes decisions based on local observation $o_{k}(t)$. To achieve topology-invariant learning, absolute coordinates are transformed into a UAV-centric relative frame:
$o_{k}(t)=\left[\mathbf{x}_{k}^{\text{self}},\mathbf{x}_{k}^{\text{GBS}},\mathbf{X}_{k}^{\text{neigh}},\mathbf{x}_{k}^{\text{user}}\right],$ (17)
where components capture specific environmental features:
@@
  • Self-State ($\mathbf{x}_{k}^{\text{self}}$): Kinematic information of the agent, defined as $\mathbf{x}_{k}^{\text{self}}=[z_{k}(t),\mathbf{v}_{k}(t)]$. Altitude $z_{k}(t)$ and velocity $\mathbf{v}_{k}(t)$ are normalized by the safe corridor and $V_{\max}^{\text{UAV}}$.
@@
  • GBS Anchor ($\mathbf{x}_{k}^{\text{GBS}}$): Relative displacement to the fixed ground station, expressed as $\mathbf{x}_{k}^{\text{GBS}}=\mathbf{p}_{\text{GBS}}-\mathbf{p}_{k}(t)$. This allows the swarm to maintain a consistent coordinate origin after other nodes fail.
@@
  • Neighborhood Topology ($\mathbf{X}_{k}^{\text{neigh}}$): Relative features from neighboring UAVs within $R{\text{comm}}$. For each neighbor $j\in\mathcal{V}_{U}(t)\setminus\{k\}$, features include relative position $\Delta\mathbf{p}_{kj}(t)=\mathbf{p}_{j}(t)-\mathbf{p}_{k}(t)$ and velocity $\mathbf{v}_{j}(t)$. The set $\mathbf{X}_{k}^{\text{neigh}}=\{[\Delta\mathbf{p}_{kj}(t),\mathbf{v}_{j}(t)]\}_{j\in\mathcal{V}_{U}(t)\setminus\{k\}}$ is processed by attention to re-weight coordination following failures.
@@
  • User Distribution ($\mathbf{x}_{k}^{\text{user}}$): Perceived as localized user densities or aggregated centroids within the footprint. They provide spatial cues for tracking high-mobility platoons and identifying service demand.
IV-C3 Action Space $\mathcal{A}$
@@
The discrete action space $\mathcal{A}$ maps $a_{k}(t)\in\mathcal{A}$ to specific target velocity vectors. To simulate inertia and ensure smooth motion, the velocity $\mathbf{v}_{k}(t+1)$ is updated as:
$\mathbf{v}_{k}(t+1)=\beta\mathbf{v}_{k}(t)+(1-\beta)\mathbf{v}_{\text{target}}(a_{k}(t))+\bm{\epsilon},$ (18)
@@
where $\beta\in[0,1)$ is the inertia coefficient and $\bm{\epsilon}$ is mechanical noise. Resulting trajectories remain continuous, and the target velocity norm is scaled to satisfy $\|\mathbf{v}_{k}(t+1)\|\leq V_{\max}^{\text{UAV}}$ for all operational UAVs in $\mathcal{V}_{U}(t)$.
IV-C4 Cooperative Reward Function $\mathcal{R}$
Agents share a common reward signal $r(t)=U(t)$ based on the utility in Section IV-A6. The primary objective is to find an optimal joint policy $\bm{\pi}$ that maximizes the expected discounted return:
$J(\bm{\pi})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(t)\right],$ (19)
@@ IV-C4 Cooperative Reward Function $\mathcal{R}$
where $\gamma\in[0,1)$ is the discount factor. Maximizing this return incentivizes surviving UAVs to autonomously re-configure their spatial topology to fill service gaps while maintaining stability. This collective reward mechanism ensures individual actions align with global performance objectives under node failure.
@@
V Topology-Aware Graph-based Multi-Agent Reinforcement Learning (TAG-MAPPO)
@@
In this section, we propose the Topology-Aware Graph-based Multi-Agent Reinforcement Learning (TAG-MAPPO) framework. This framework is specifically engineered to surmount the challenges of high-mobility tracking and rapid topology deformation caused by both user movement and sudden node failures. To address the performance degradation typical of existing solutions in volatile environments, TAG-MAPPO adopts a Centralized Training with Decentralized Execution (CTDE) paradigm augmented by a specialized attention mechanism. Recent literature confirms that graph-based learning is uniquely suited for such dynamic wireless resource management tasks because of its inherent scalability and generalization capabilities [7].
@@
V-A Overview of the TAG-MAPPO Architecture
The overall architectural flow of the proposed TAG-MAPPO framework is illustrated in Fig. 2. The processing pipeline is strictly structured to differentiate between the online execution phase and the offline centralized training phase.
Figure 2: The schematic architecture of the proposed TAG-MAPPO framework: (1) Decentralized Execution (Actor): Each UAV agent utilizes a lightweight, shared MLP encoder to map local observations $o_{k}(t)$ to actions, ensuring real-time responsiveness. (2) Centralized Training (Critic): The critic leverages a Topology-Aware Graph Attention (TA-GAT) mechanism to estimate state values $V_{\phi}(s_{t})$. This module employs a decoupled dual-path strategy. A Random Observation Shuffling (ROS) operator is applied to the neighbor feature set $\{\mathbf{h}_{j}\}_{j\in\mathcal{N}_{k}(t)}$ to ensure permutation invariance. Simultaneously, a bold red skip connection preserves the ego-state information $\mathbf{h}_{k}$ to maintain operational consistency. This architecture enables the framework to perform precise relational reasoning and rapid spatial reconfiguration, which ensures robust coordination even under sudden node failures.